# Your Latent Reasoning is Secretly Policy Improvement Operator

### v3

@@ 6 Conclusion
We demonstrate that small, iterative reasoning models can achieve competitive performance on complex reasoning tasks such as the Abstraction and Reasoning Corpus, challenging the dominance of large-scale language models. By reinterpreting TRMs through the lens of reinforcement learning, we reveal that TRMs implicitly perform policy improvement, where a latent ”working memory” state guides the model toward better solutions over recursive steps.
The key contribution is Deep Improvement Supervision, builds on this insight by introducing a structured, stepwise training regime. DIS provides intermediate targets through a discrete diffusion process, transforming the challenging problem of long-term credit assignment into a more tractable supervised learning task (Ho and Salimans, 2022; Frans et al., 2025).
Our approach not only simplifies training by eliminating the need for learned halting mechanisms but also enhances efficiency, reducing the number of forward passes by 18x with high accuracy.
@@ Appendix B Proofs
$A_{s}(a)\;=\;\big(\ell_{c}^{s}[a]-\ell_{u}^{s}[a]\big)\;-\;\Big(\log Z_{c}^{s}-\log Z_{u}^{s}\Big)\;=\;\Delta\ell^{s}(a)\;-\;C_{s},$
where $Z_{c}^{s}=\sum_{a^{\prime}}e^{\ell_{c}^{s}[a^{\prime}]}$, $Z_{u}^{s}=\sum_{a^{\prime}}e^{\ell_{u}^{s}[a^{\prime}]}$, and $C_{s}$ does not depend on $a$. Therefore, for any distribution $\mu$ over actions,

### v4

@@ 6 Conclusion
We demonstrate that small iterative reasoning models can achieve competitive performance on complex reasoning tasks such as the ARC. By reinterpreting TRMs through the lens of reinforcement learning, we reveal that TRMs implicitly perform policy improvement, where a latent working memory state guides the model toward better solutions over recursive steps. The key contribution is Deep Improvement Supervision, builds on this insight by introducing a structured, stepwise training regime. DIS provides intermediate targets through a discrete diffusion process (Ho and Salimans, 2022; Frans et al., 2025), transforming the challenging problem of long-term credit assignment into a more tractable supervised learning task. Our approach not only simplifies training by eliminating the need for learned halting mechanisms but also enhances efficiency, reducing the number of forward passes by 18x with high accuracy.
@@ Appendix B Proofs
$A_{s}(a)\;=\;\big(\ell_{c}^{s}[a]-\ell_{u}^{s}[a]\big)\;-\;\Big(\log\mathbf{z}_{c}^{s}-\log\mathbf{z}_{u}^{s}\Big)\;=\;\Delta\ell^{s}(a)\;-\;C_{s},$
where $\mathbf{z}_{c}^{s}=\sum_{a^{\prime}}e^{\ell_{c}^{s}[a^{\prime}]}$, $\mathbf{z}_{u}^{s}=\sum_{a^{\prime}}e^{\ell_{u}^{s}[a^{\prime}]}$, and $C_{s}$ does not depend on $a$. Therefore, for any distribution $\mu$ over actions,